# Classification of abelian Schur groups I

### v1

@@ 7.3. Dense $S$-rings
One can see that
$\underline{X}^{2}=\sum\limits_{h\in X_{h}}\underline{X_{h}}^{2}+\sum\limits_{h,h^{\prime}\in X_{H}}hh^{\prime}\underline{X_{h}}\cdot\underline{X_{h^{\prime}}}\in\mathcal{A}.$
Since $D$ is an $\mathcal{A}$-subgroup, we conclude that
$\xi=\sum\limits_{h\in X_{h}}\underline{X_{h}}^{2}\in\mathcal{A}.$
Due to Eqs. (1), (22), and (23),
$\xi=\sum\limits_{h\in X_{h}}(\underline{X_{h}\setminus X_{h,p}}+\underline{X_{h,p}})^{2}=p\eta+2(p-1)/|X_{H}|\sum\limits_{h\in X_{H}}(\underline{X_{h}}-\underline{X_{h,p}})+\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}$
for some $\eta\in\mathbb{Z}D$. The above equality implies that every element from $D\setminus P$ enters $\xi$ with the coefficient equal to $2(p-1)/|X_{H}|$ modulo $p$. On the other hand, the above expression for $\xi$ and Eqs. (24) and (25) imply that every element from $P^{\#}$ enters $\xi$ with coefficient equal to $(p-1)/|X_{H}|$ or $(p-1)/|X_{H}|-1$ modulo $p$. However, each element from $D^{\#}$ must enter $\xi$ with the same coefficient because $D^{\#}\in\mathcal{S}(\mathcal{A})$, a contradiction.
Case 2: $\mathcal{A}_{D}$ is cyclotomic. In this case, all basic sets of $\mathcal{A}_{D}$ are regular. Let $K_{D}\leq\operatorname{Aut}(D)$ such that $\mathcal{A}_{D}=\operatorname{Cyc}(K_{D},D)$. We start the proof in this case with the following claim.
@@
It is enough to prove the claim for every prime $r$. Indeed, if $r$ is composite, then successively applying the claim for each prime divisor of $r$, we obtain the required. If $r$ is coprime to $2n$, then the claim follows from Lemma 2.3. Lemma 3.2(1) implies that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$. If $r=2$, then $X^{(r)}=X^{(2)}=X_{D}^{(2)}\in\mathcal{S}(\mathcal{A}_{D})$ by Lemma 2.3. Further, we assume that $r$ is an odd prime divisor of $n$.
Put $Y=X^{(r)}$. Let $R$ be the (unique) subgroup of $D$ of order $R$. Since $\mathcal{A}_{D}$ is cyclotomic, $R$ is an $\mathcal{A}$-subgroup. Observe that $R\nleq\operatorname{rad}(X_{D})$. Indeed, if $n=p^{k}$, then this follows from Lemma 6.8(2), whereas if $n=pq$, then this immediately follows from the the fact that $\mathcal{A}$ is cyclotomic with trivial radical. Therefore $R\nleq\operatorname{rad}(X)$ and hence $Y=X^{[r]}$. Lemma 2.4 yields that $Y$ is an $\mathcal{A}$-set.
@@
Due to Lemma 3.2(1), the set $Y_{D}$ is an $\mathcal{A}_{D}$-set. One can see that $Y_{D}=X_{D}^{(r)}$. If $n=p^{k}$, then $|Y_{D}|=|X_{D}|$ because all the elements of $X_{D}$ lie in pairwise distinct $R$-cosets (Lemma 6.8(1)). From Lemma 6.8(2) it follows that all nontrivial basic set of $\mathcal{A}_{D}$ has the same size and hence Eq. (26) holds. If $n=pq$, then $R$ and the Hall $r^{\prime}$-subgroup $U$ of $D$ are $\mathcal{A}$-subgroups because $\mathcal{A}_{D}$ is cyclotomic. So $H$, $U$, and $R$ are $\mathcal{A}$-subgroups such that $G=H\times U\times R$. As $R\nleq\operatorname{rad}(X)$, one can see that $Y=X_{H\times U}$ and consequently $Y_{D}=(X_{H\times U})_{D}=X_{U}$. Therefore $Y_{D}\in\mathcal{S}(\mathcal{A}_{U})$ by Lemma 3.2(1) applied to $G_{1}=H\times R$ and $G_{2}=U$. Thus, Eq. (26) holds.
Assume that $Y\notin\mathcal{S}(\mathcal{A})$. Then there exists $Z\in\mathcal{S}(\mathcal{A})$ such that $Z\subsetneq Y$. Suppose that $\lambda_{X}=1$. Then $\lambda_{Y}=1$ and hence $|Y|=|Y_{D}|$. Together with $Z\subsetneq Y$, this implies that $Z_{D}\subsetneq Y_{D}$. Due to Eq. (26), we conclude that $Z_{D}\notin\mathcal{S}(\mathcal{A}_{D})$, a contradiction to Lemma 3.2(1).
Now suppose that $\lambda_{X}\neq 1$. Then $X=X_{H}\times X_{D}$ by Lemma 7.10 and $\lambda_{X}=|X_{H}|\geq 2$. So $Y=Y_{H}\times Y_{D}=X_{H}\times Y_{D}$. As $Z_{H},Y_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $Z_{H}\subseteq Y_{H}$, we have $Z_{H}=Y_{H}$. Besides, $Z_{D}=Y_{D}$ by Eq. (26). Since $Z\subsetneq Y$, we have $\lambda_{Z}=1$ by Lemma 7.10.
@@
In view of Lemma 3.2(1) and the assumptions on $\mathcal{A}_{H}$ and $\mathcal{A}_{D}$, one can choose a basic set $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that
@@
and $X_{D}$ consists of generators of $D$. Clearly, $X_{H}\in\operatorname{Orb}(K_{H},H)$ and $X_{D}\in\operatorname{Orb}(K_{D},D)$. Due to Lemma 7.10, $\lambda_{X}\in\{1,|X_{H}|\}$. Let $\mathcal{S}_{0}=\{X^{(r)}:~r\in\mathbb{Z}\}$. Claim $1$ implies that $\mathcal{S}_{0}\subseteq\mathcal{S}(\mathcal{A})$.
@@
If $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D})\cup\mathcal{S}_{0})$, then $|Y_{H}|=1$ and, in particular, $Y=Y_{H}\times Y_{D}$.
@@
If $Y_{H}=X_{H}$, then $Y\subseteq H$ or $Y\in\mathcal{S}_{0}$, a contradiction to the choice of $Y$. So $Y_{H}\neq X_{H}$. Since $X_{H},Y_{H}\in\mathcal{S}(\mathcal{A}_{H})$ (Lemma 3.2(1)) and $|X_{H}|>1$, we obtain $|Y_{H}|=1$ as required. ∎
If $\lambda_{X}=|X_{H}|$, then $X=X_{H}\times X_{D}$. Let $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Then $Y=Y_{H}\times Y_{D}$. Indeed, this is clear if $Y\in\mathcal{S}_{0}$ and follows from Claim $2$ otherwise. Therefore $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, a contradiction to the assumption of the proposition. Thus,
@@
Lemma 2.5 implies that $X_{h}$ is an orbit of some $K^{0}_{D}\leq\operatorname{Aut}(D)$ for every $h\in H$. Since both $K_{D}$ and $K^{0}_{D}$ are semiregular on $X_{D}$, we conclude that $K_{0}\leq K_{D}$ and $|K_{D}:K^{0}_{D}|=|X_{H}|$. As $|K_{H}|=|X_{H}|=|K_{D}/K^{0}_{D}|=|K_{H}|\leq 3$, there is an isomorphism $\psi$ from $K_{H}$ to $K_{D}/K^{0}_{D}$ such that
$(X_{h})^{\sigma^{\psi}}=X_{h^{\sigma}}.$
Put
@@
By the definition, $K$ is a subdirect product of $K_{H}$ and $K_{D}$.
Put $\mathcal{B}=\operatorname{Cyc}(K,G)$. To prove that $\mathcal{A}$ is cyclotomic, it is enough to show that $\mathcal{A}=\mathcal{B}$. One can see that $\mathcal{A}_{H}=\mathcal{B}_{H}$ ($\mathcal{A}_{D}=\mathcal{B}_{D}$, respectively) because $K^{H}=K_{H}$ ($K^{D}=K_{D}$, respectively). As $X_{H}$ and $X_{D}$ are orbits of $K_{H}$ and $K_{D}$, respectively, we conclude that $X\in\operatorname{Orb}(K,G)$ and hence $X\in\mathcal{S}(\mathcal{B})$. Due to Claim $1$ applied to $\mathcal{A}$ and to $\mathcal{B}$, we have
$\mathcal{S}_{0}\subseteq\mathcal{S}(\mathcal{A})\cap\mathcal{S}(\mathcal{B}).$
We are done if $\mathcal{S}(\mathcal{A})=\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D})\cup\mathcal{S}_{0}$. Otherwise, $Y=Y_{H}\times Y_{D}$ for every $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D})\cup\mathcal{S}_{0})$ by Claim $2$ and hence $Y\in\operatorname{Orb}(K,G)=\mathcal{S}(\mathcal{B})$ by the definition of $K$. Thus, $\mathcal{A}=\mathcal{B}$ and we are done.
Now let us prove that $\mathcal{A}$ is $2$-minimal. Let $x\in X$. It is enough to show that the pointwise stabilizer $\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial. Indeed, then $X$ is a faithful regular orbit of $\operatorname{Aut}(\mathcal{A})_{e}$ and $\mathcal{A}$ is $2$-minimal by [21, Lemma 8.2]. It suffices to verify that every $f\in\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial.
@@
Since $\mathcal{A}=\operatorname{Cyc}(K,G)$ and $\mathcal{A}$ is $2$-minimal, we conclude that $\operatorname{Aut}(\mathcal{A})=G_{r}\rtimes K$. This implies that $\mathcal{A}$ is normal and $|\operatorname{Aut}(\mathcal{A})^{H}|=|H||K^{H}|\leq 12$ which finishes the proof of the proposition. ∎
Proposition 7.12.
Suppose that $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, where $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $U$ is an $\mathcal{A}_{D}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Then one of the following statements holds:
@@
At first, suppose that $n=pq$. Without loss of generality, we may assume that $|U|=|L|=p$. Then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$. So $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product by Lemma 5.3 applied to $H_{1}=H$ and $P_{1}=L$ and Statement $(1)$ of the proposition holds.
Now suppose that $n=p^{k}$.
@@
If $L\nleq\operatorname{rad}(X)$ for some $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$, then $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, $L=P$, and $X$ is one of the following forms
(27) $h_{1}\{x\}\cup h_{2}\{xx_{0}\}\cup h_{3}\{xx_{0}^{2}\},~h_{1}\{x,x^{-1}\}\cup h_{2}\{xx_{0},x^{-1}x_{0}^{2}\}\cup h_{3}\{xx_{0}^{2},x^{-1}x_{0}\},$
where $h_{1},h_{2},h_{3}$ are pairwise distinct nontrivial elements of $H$, $x_{0}$ is a generator of $P$, and $x\in D$. In particular, $|X|=|X_{D}|\leq 6$.
@@
Note that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$ by Lemma 3.2(1) and hence $L\leq\operatorname{rad}(X_{D})$. If $X=X_{H}\times X_{D}$, then $L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. Otherwise, $|X_{H}|\in\{2,3\}$ by Lemma 7.8 and $\lambda_{X}=1$ by Lemma 7.10. The number $\mu_{X}=|X\cap Lx|$ does not depend on $x\in X$ by Lemma 2.1. Since $\lambda_{X}=1$, we conclude that $\mu_{X}=|L|$ or $\mu_{X}=|L|/|X_{H}|$. In the former case, $L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. In the latter one, since $|X_{H}|\in\{2,3\}$ and $|L|$ is a $p$-power for an odd prime $p$, we obtain $|X_{H}|=p=3$. This implies that $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $X_{H}=H^{\#}$.
Let $X_{D}$ be regular. Then all $X_{h}$’s, $h\in H^{\#}$, are orbits of some $K\leq\operatorname{Aut}(\langle X_{D}\rangle)$ by Lemma 2.5. Since $L\nleq\operatorname{rad}(X)$, we obtain $P\nleq\operatorname{rad}(X_{h})$ and hence $|\operatorname{rad}(X_{h})|=1$ for every $h\in H^{\#}$. Lemma 6.8(1) implies that $|K|=|X_{h}|\leq 2$ for every $h\in H^{\#}$. Therefore each $X_{h}$ is singleton or each $X_{h}$ is of the form $\{x,x^{-1}\}$, $x\in D$. Due to $\lambda_{X}=1$, all $X_{h}$’s are pairwise disjoint and consequently
@@
As $L\leq\operatorname{rad}(X_{D})$, we have $L=\operatorname{rad}(X_{D})=P$ and hence $X_{D}=Px$ or $X_{D}=Px\cup Px^{-1}$. Together with the previous discussion, this yields that $X$ is one of the required forms.
Now let $X_{D}$ be nonregular. Then $X_{D}=V\setminus N$ for some $\mathcal{A}_{D}$-subgroups $V>N\geq U$ by Lemma 6.11. Let $m$ be the minimum of orders of the elements from $X_{D}$. If $P\nleq\operatorname{rad}(X_{h,l})$ for some $h\in H^{\#}$ and $l>m$, then

### v3

@@ 7.3. Dense $S$-rings
Put
$\xi=\sum\limits_{h\in X_{H}}\underline{X_{h}}^{2}\in\mathbb{Z}D.$
One can compute directly that
$\underline{X}^{2}=\xi+\sum\limits_{h,h^{\prime}\in X_{H}}hh^{\prime}\underline{X_{h}}\cdot\underline{X_{h^{\prime}}}\in\mathcal{A}.$
Observe that only elements from $G\setminus D$ enter the second sum on the right-hand side of the above equality. Since $D$ is an $\mathcal{A}$-subgroup, we conclude that $\xi\in\mathcal{A}$. Further, one can compute $\xi$ as follows:
$\xi=\sum\limits_{h\in X_{H}}(\underline{X_{h}\setminus X_{h,p}}+\underline{X_{h,p}})^{2}=\sum\limits_{h\in X_{H}}(\underline{X_{h}\setminus X_{h,p}})^{2}+2\sum\limits_{h\in X_{H}}\underline{X_{h}\setminus X_{h,p}}\cdot\underline{X_{h,p}}+\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}.$
Each element appearing in the first of three last sums enters with a coefficient divisible by $p$ due to Eq. (24) and hence
$\sum\limits_{h\in X_{H}}(\underline{X_{h}\setminus X_{h,p}})^{2}=p\eta$
for some $\eta\in\mathbb{Z}D$. Recall that $X_{h,p}\subseteq P\leq\operatorname{rad}(X_{h}\setminus X_{h,p})$ for every $h\in X_{H}$, where the latter inequality holds by Eq. (24). Therefore
$\sum\limits_{h\in X_{H}}\underline{X_{h}\setminus X_{h,p}}\cdot\underline{X_{h,p}}=\sum\limits_{h\in X_{H}}|X_{h,p}|\underline{X_{h}\setminus X_{h,p}}=2(p-1)/|X_{H}|\sum\limits_{h\in X_{H}}(\underline{X_{h}}-\underline{X_{h,p}}),$
where the first equality follows from Eq. (1), whereas the second one from Eq. (25). Thus,
$\xi=p\eta+2(p-1)/|X_{H}|\sum\limits_{h\in X_{H}}(\underline{X_{h}}-\underline{X_{h,p}})+\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}.$
The above expression for $\xi$ implies that every element from $D\setminus P$ enters $\xi$ with the coefficient equal to $2(p-1)/|X_{H}|$ modulo $p$. On the other hand, this expression and Eqs. (26) and (27) imply that every element from $P^{\#}$ enters $\xi$ with coefficient equal to $(p-1)/|X_{H}|$ or $(p-1)/|X_{H}|-1$ modulo $p$. It is easy to see that none of the above numbers is equal to $2(p-1)/|X_{H}|$ modulo $p$. However, each element from $D^{\#}$ must enter $\xi$ with the same coefficient because $D^{\#}\in\mathcal{S}(\mathcal{A})$, a contradiction.
Case 2: $\mathcal{A}_{D}$ is cyclotomic. In this case, all basic sets of $\mathcal{A}_{D}$ are regular. Let $K_{D}\leq\operatorname{Aut}(D)$ be such that $\mathcal{A}_{D}=\operatorname{Cyc}(K_{D},D)$.
@@
Without loss of generality, we may assume that $r$ is prime. If $r$ is coprime to $2n$, then the claim follows from Lemma 2.3. Lemma 3.2(1) implies that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$. If $r=2$, then $X^{(r)}=X^{(2)}=X_{D}^{(2)}\in\mathcal{S}(\mathcal{A}_{D})$ by Lemma 2.3.
It remains to consider the case when $r$ is an odd prime divisor of $n$. Put $Y=X^{(r)}$. Let $R$ be the (unique) subgroup of $D$ of order $r$. Since $\mathcal{A}_{D}$ is cyclotomic, $R$ is an $\mathcal{A}$-subgroup. Observe that $R\nleq\operatorname{rad}(X_{D})$. Indeed, if $n=p^{k}$, then this follows from Lemma 6.8(2), whereas if $n=pq$, then this immediately follows from the the fact that $\mathcal{A}$ is cyclotomic with trivial radical. Therefore $R\nleq\operatorname{rad}(X)$ and hence $Y=X^{[r]}$. Lemma 2.4 yields that $Y$ is an $\mathcal{A}$-set.
@@
One can see that $Y_{D}=X_{D}^{(r)}$. If $n=p^{k}$, then $|Y_{D}|=|X_{D}|$ because all the elements of $X_{D}$ lie in pairwise distinct $R$-cosets (Lemma 6.8(1)). From Lemma 6.8(2) it follows that all nontrivial basic set of $\mathcal{A}_{D}$ has the same size and hence Eq. (28) holds. If $n=pq$, then $R$ and the Hall $r^{\prime}$-subgroup $U$ of $D$ are $\mathcal{A}$-subgroups because $\mathcal{A}_{D}$ is cyclotomic. So $H$, $U$, and $R$ are $\mathcal{A}$-subgroups such that $G=H\times U\times R$. As $R\nleq\operatorname{rad}(X)$, one can see that $Y=X_{H\times U}$ and consequently $Y_{D}=(X_{H\times U})_{D}=X_{U}$. Therefore $Y_{D}\in\mathcal{S}(\mathcal{A}_{U})$ by Lemma 3.2(1) applied to $G_{1}=H\times R$ and $G_{2}=U$. Thus, Eq. (28) holds.
To complete the proof, let us show that $Y\in\mathcal{S}(\mathcal{A})$. Assume the contrary that $Y\notin\mathcal{S}(\mathcal{A})$. Then there exists $Z\in\mathcal{S}(\mathcal{A})$ such that $Z\subsetneq Y$. Recall that $\lambda_{X}=|Hx\cap X|$ does not depend on $x\in X$. We have $\lambda_{X}\neq 1$. Indeed, if $\lambda_{X}=1$, then $\lambda_{Y}=1$ and hence $|Y|=|Y_{D}|$. Together with $Z\subsetneq Y$, this implies that $Z_{D}\subsetneq Y_{D}$. Due to Eq. (28), we conclude that $Z_{D}\notin\mathcal{S}(\mathcal{A}_{D})$, a contradiction to Lemma 3.2(1).
Since $\lambda_{X}\neq 1$, we have $X=X_{H}\times X_{D}$ by Lemma 7.11 and $\lambda_{X}=|X_{H}|\geq 2$. So $Y=Y_{H}\times Y_{D}=X_{H}\times Y_{D}$. As $Z_{H},Y_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $Z_{H}\subseteq Y_{H}$, we have $Z_{H}=Y_{H}$. Besides, $Z_{D}=Y_{D}$ by Eq. (28). Since $Z\subsetneq Y$, we have $\lambda_{Z}=1$ by Lemma 7.11.
@@
Since $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order $2$ or $3$, and $\mathcal{A}_{D}=\operatorname{Cyc}(K_{D},D)$, one can choose a basic set $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that
@@
and $X_{D}$ consists of generators of $D$. Clearly, $X_{H}\in\operatorname{Orb}(K_{H},H)$ and $X_{D}\in\operatorname{Orb}(K_{D},D)$.
@@
Let $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D}))$. Then $Y=X^{(r)}$ for some $r\in\mathbb{Z}$ or $|Y_{H}|=1$ and $Y=Y_{H}\times Y_{D}$.
@@
Let $Y_{H}=X_{H}$. Since $Y\notin\mathcal{S}(\mathcal{A}_{H})$ by the condition of the claim and $X_{D}$ consists of generators of $D$, we conclude that $Y\cap X^{(r)}\neq\varnothing$ for some $r\in\mathbb{Z}$. Claim $1$ implies that $X^{(r)}\in\mathcal{S}(\mathcal{A})$ and hence $Y=X^{(r)}$ as required. Now let $Y_{H}\neq X_{H}$. Since $X_{H},Y_{H}\subseteq H^{\#}$ and $|X_{H}|>1$, we obtain $|Y_{H}|=1$ and hence $Y=Y_{H}\times Y_{D}$ as required. ∎
If $X=X_{H}\times X_{D}$, then $Y=Y_{H}\times Y_{D}$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ by Claim $2$ and hence $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ which contradicts to the supposition of the theorem. Therefore $X\neq X_{H}\times X_{D}$. Together with Lemma 7.11, this implies that
@@
Further, we are going to construct $K\leq\operatorname{Aut}(G)$ such that
(29) $X\in\operatorname{Orb}(K,G).$
Since $X_{D}\in\operatorname{Orb}(K_{D},D)$, the set $X_{D}$ is regular. So $X_{h}=X_{h,p^{k}}$ for every $h\in X_{H}$. Lemma 2.5 implies that $X_{h}$ is an orbit of some $K^{0}_{D}\leq\operatorname{Aut}(D)$ for every $h\in H$. Clearly, $K^{0}_{D}\leq K_{D}$, $K_{D}$ acts transitively on the set $\{X_{h}:~h\in X_{H}\}=\operatorname{Orb}(K^{0}_{D},X_{D})$, and $K^{0}_{D}$ is the kernel of this action. As $\lambda_{X}=1$, we have $|K_{D}:K^{0}_{D}|=|X_{H}|=|K_{H}|\leq 3$. Therefore $K_{H}\cong K_{D}/K^{0}_{D}\cong C_{|X_{H}|}$. Let $\sigma_{0}$ be a generator of $K_{H}$. Since $|K_{D}:K^{0}_{D}|=|K_{H}|\leq 3$, there is $\tau_{0}\in K_{D}$ such that
(30) $X_{h}^{\tau_{0}}=X_{h^{\sigma_{0}}}$
for every $h\in X_{H}$. Let $\psi$ be an isomorphism from $K_{H}$ to $K_{D}/K^{0}_{D}$ such that $\sigma_{0}^{\psi}=K^{0}_{D}\tau_{0}$ and
@@
Eq. (30) yields that $X$ is $K$-invariant. As $K_{H}$ and $K_{D}$ are transitive on $X_{H}$ and $X_{D}$, respectively, so is $K$ on $X$. Thus, Eq. (29) holds.
To prove that $\mathcal{A}$ is cyclotomic, it is enough to show that $\mathcal{A}=\operatorname{Cyc}(K,G)$. One can see that $\mathcal{A}_{H}=\operatorname{Cyc}(K,G)_{H}$ ($\mathcal{A}_{D}=\operatorname{Cyc}(K,G)_{D}$, respectively) because $K^{H}=K_{H}$ ($K^{D}=K_{D}$, respectively). Let $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D}))$. Then $Y=X^{(r)}$ for some $r\in\mathbb{Z}$ or $|Y_{H}|=1$ and $Y=Y_{H}\times Y_{D}$ by Claim $2$. In the former case, $Y\in\operatorname{Orb}(K,G)$ by Eq. (29), whereas in the latter one, by the definition of $K$. Thus, every basic set of $\mathcal{A}$ is an orbit of $K$, i.e. $\mathcal{A}=\operatorname{Cyc}(K,G)$ as desired.
Now let us prove that $\mathcal{A}$ is $2$-minimal. Let $x\in X$. It is enough to show that the pointwise stabilizer $\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial. Indeed, then $X$ is a faithful regular orbit of $\operatorname{Aut}(\mathcal{A})_{e}$ and $\mathcal{A}$ is $2$-minimal by [23, Lemma 8.2]. It suffices to verify that every $f\in\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial.
@@
Since $\mathcal{A}=\operatorname{Cyc}(K,G)$ and $\mathcal{A}$ is $2$-minimal, we have $\operatorname{Aut}(\mathcal{A})=G_{r}\rtimes K$. This implies that $\mathcal{A}$ is normal and $|\operatorname{Aut}(\mathcal{A})^{H}|=|H||K^{H}|\leq 12$ which completes the proof of the theorem. ∎
Theorem 7.13.
Let $\mathcal{A}$ be dense. Suppose that $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, where $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $U$ is an $\mathcal{A}_{D}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Then one of the following statements holds:
@@
At first, suppose that $n=pq$. As $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, we may assume without loss of generality that $|U|=|L|=p$. Then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$. So $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product by Lemma 5.3 applied to $H_{1}=H$ and $P_{1}=L$ and Statement $(1)$ of the theorem holds as desired.
Now suppose that $n=p^{k}$. Since $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, we have $k\geq 2$.
@@
If there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$ such that $L\nleq\operatorname{rad}(X)$, then $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, $L=P$, and every such $X$ is equal to
(31) $h_{1}\{x\}\cup h_{2}\{xx_{0}\}\cup h_{3}\{xx_{0}^{2}\}~\text{or}~h_{1}\{x,x^{-1}\}\cup h_{2}\{xx_{0},x^{-1}x_{0}^{2}\}\cup h_{3}\{xx_{0}^{2},x^{-1}x_{0}\},$
where $h_{1},h_{2},h_{3}$ are pairwise distinct nonidentity elements of $H$, $x_{0}$ is a generator of $P$, and $x\in D$. In particular, $|X|=|X_{D}|\leq 6$.
@@
We have $X\neq X_{H}\times X_{D}$ because otherwise $L\leq\operatorname{rad}(X_{D})$ and hence $L\leq\operatorname{rad}(X)$ that contradicts to the assumption of the lemma. Hence $|X_{H}|\in\{2,3\}$ by Lemma 7.9 and
$\lambda_{X}=|X\cap Hx|=1~\text{for every}~x\in X$
by Lemma 7.11. The number $\mu_{X}=|X\cap Lx|$ does not depend on $x\in X$ by Lemma 2.1. Since $\lambda_{X}=1$, we conclude that $\mu_{X}=|L|$ or $\mu_{X}=|L|/|X_{H}|$. In the former case, $L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. In the latter one, since $|X_{H}|\in\{2,3\}$ and $|L|$ is a $p$-power for an odd prime $p$, we obtain $|X_{H}|=p=3$. This implies that $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $X_{H}=H^{\#}$.
Let $X_{D}$ be regular. Then all the $X_{h}$, $h\in H^{\#}$, are orbits of some $K\leq\operatorname{Aut}(\langle X_{D}\rangle)$ by Lemma 2.5. Since $L\nleq\operatorname{rad}(X)$, we obtain $P\nleq\operatorname{rad}(X_{h})$ and hence $|\operatorname{rad}(X_{h})|=1$ for every $h\in H^{\#}$. Lemma 6.8(1) implies that $|K|=|X_{h}|\leq 2$ for every $h\in H^{\#}$. Therefore each $X_{h}$ is a singleton or each $X_{h}$ is of the form $\{x,x^{-1}\}$, $x\in D$. Due to $\lambda_{X}=1$, all the $X_{h}$ are pairwise disjoint and consequently
@@
As $L\leq\operatorname{rad}(X_{D})$, we have $L=\operatorname{rad}(X_{D})=P$ and hence $X_{D}=Px$ or $X_{D}=Px\cup Px^{-1}$ for some $x\in D$. In the former case, each $X_{h}$ is a singleton and hence $X$ is of the first form from Eq. (31), whereas in the latter one, each $X_{h}$ consists of two mutually inverse elements and hence $X$ is of the second form from Eq. (31) as desired.
Now let $X_{D}$ be nonregular. Then $X_{D}=V\setminus N$ for some $\mathcal{A}_{D}$-subgroups $V>N\geq U$ such that $|V:N|\geq 9$ by Lemma 6.11. Let $m$ be the minimum of orders of the elements from $X_{D}$. If $P\nleq\operatorname{rad}(X_{h,l})$ for some $h\in H^{\#}$ and $l>m$, then